# HyperspectralViTs: General Hyperspectral Models for On-board Remote Sensing

### v2

@@ Abstract
On-board processing of hyperspectral data with machine learning models would enable unprecedented amount of autonomy for a wide range of tasks, for example methane detection or mineral identification. This can enable early warning system and could allow new capabilities such as automated scheduling across constellations of satellites. Classical methods suffer from high false positive rates and previous deep learning models exhibit prohibitive computational requirements. We propose fast and accurate machine learning architectures which support end-to-end training with data of high spectral dimension without relying on hand-crafted products or spectral band compression preprocessing. We evaluate our models on two tasks related to hyperspectral data processing. With our proposed general architectures, we improve the F1 score of the previous methane detection state-of-the-art models by 27% on a newly created synthetic dataset and by 13% on the previously released large benchmark dataset. We also demonstrate that training models on the synthetic dataset improves performance of models finetuned on the dataset of real events by 6.9% in F1 score in contrast with training from scratch. On a newly created dataset for mineral identification, our models provide 3.5% improvement in the F1 score in contrast to the default versions of the models. With our proposed models we improve the inference speed by 85% in contrast to previous classical and deep learning approaches by removing the dependency on classically computed features. With our architecture, one capture from the EMIT sensor can be processed within 30 seconds on realistic proxy of the ION-SCV 004 satellite.
@@
publicationid: pubid: $0$
@@ I Introduction
Mineral identification is a task related to scene decomposition into individual endmembers from a spectral library. On its own it could serve as a strong general task highly relevant for future Foundation models working with hyperspectral data. We note that one of the example downstream tasks of interest could also be methane detection, as it has been shown that having information about scene background is beneficial for separation of methane leak events from confounders [4].
@@
Figure 1: Illustration of the limitations of previous approaches and our proposed HyperspectralViTs, adapted Transformer based architecture.
@@ I-A Literature review
As has been identified by [29] and several overview papers [30, 31, 32], research in machine learning for processing hyperspectral data has been limited mainly due to the lack of large, publicly available datasets with high quality annotations. Many works still use very small datasets such as Indian Pines [33], University of Pavia, Houston or Salinas, all of which contain only a few data captures. As further noted by [30], this data is often divided into almost equal train and test subsets, which leads to model overfitting and results with very high accuracies almost regardless of the used model. In the domain of methane leak detection, only the recent works of [34, 4] release processed datasets from raw data captures, [34] furthermore refines the available labels. The works of [35, 36] propose machine learning models for detection and quantification of methane leak events in hyperspectral data, however the used datasets and code/model weights are not published. In contrast, we publish all our used code and trained models and present the created datasets in a machine learning ready format.
@@ II-B Mineral aggregation
To create our new mineral dataset, we adapt the steps described in the Theoretical basis document [46] released alongside the “EMITL2BMIN.001” product [47] computed from the raw EMIT data. This data contains identifications of 381 individual components (separated into two groups). These can be aggreggated into larger mineral classes, for this we select the so-called “EMIT10” subset defined in [46]. The aggregation process is shown on Figure 4. We note that we use only the binary maps of three selected mineral classes out of the total 10. We chose the subset M = {Goethite, Hematite, Kaolinite}, as these classes were abundant in the locations sampled in our dataset. We use the released annotations which were computed using the Tetracorder method [24], and as such we can’t really use Tetracorder as a classical baseline (as it would unfairly obtain perfect scores). Therefore, we consider the created dataset as having pseudo-ground truth labels, and our results only as initial proof-of-principle. We however note, that given any manually validated dataset, our models could be easily retrained with the new data.
@@ II-C Final created datasets
Firstly, we created the “OxHyperSyntheticCH4” dataset of synthetically simulated methane leak events using the labels of the STARCOP dataset with newly downloaded near-global hyperspectral datacubes from the EMIT sensor (in L1B level of data processing). In total we keep 796 large tiles (of 512x512 px) in the training dataset, 198 in the validation dataset and 200 in the test dataset. Each subset is made from completely non-overlapping sources of data (clean EMIT datacubes and labels from the STARCOP dataset from AVIRIS-NG data). In each subset we uniformly sample the available plume sizes, so that we have similar distribution of event sizes. Half of these tiles contain simulated events while the other half is kept event free to provide negative samples. We note that these large tiles (of 512x512 px) are further tiled when training the models. When using the tile size of 64 pixels with an overlap of 32 pixels, we finally get a training dataset of 174341 samples. The total size of this data is 228 GB.
Secondly, we create the “OxHyperRealCH4” dataset of real methane leak events. In total we have 279 large tiles in the training dataset, 91 in the validation dataset and 98 in the test dataset. When further tiled, the training dataset has 54805 samples (64 pixel tiles). These have reliable ground truth labels, which we calculated from the original labels provided on the EMIT data portal and consequentially manually checked. We note that this dataset is smaller than the other ones and training a machine learning model from scratch could lead to overfitting and poor generalisation, however it can also be used for model fine-tuning and to test generalisation ability of models trained on other datasets. The total size of this data is 47 GB.
Thirdly, we create “OxHyperMinerals”, a completely new dataset for mineral identification, with labels of 3 selected mineral classes and source annotation for all 381 constituents (separated into two groups). In total we have 796 large tiles in the training dataset, 198 in the validation and finally 200 in the test dataset. When further tiled, the training dataset has 174341 samples (64 pixel tiles). The large tiles exactly match the tiles used in the synthetic methane events dataset, however they do not contain the simulated events and contain more bands. The mineral dataset contains all 285 spectral bands in between spectral range of 381-2492nm. The total size of this data is 372 GB.
Finally, we also use a variant of the STARCOP dataset from [34] with all relevant spectral bands extracted from the source data. We train our models and compare them on this benchmark dataset with the previously reported results.
@@ III-A Limitations of existing approaches
We are interested in training these architectures in an end-to-end manner on the raw hyperspectral datacubes. In the methane detection task, this allows us to not depend on the classical matched filter product, which can be slow and which contains false detections of confounder materials. Models designed with typical computer vision data in mind (with one to three bands) however suffer from what we call informational bottleneck when they are used with data with large spectral dimensionality. As an example, if we select 86 bands of data which are relevant for our task of methane detection, the first layer in the vanilla versions of the used models reduces the data volume by $-95.34\%$ in the case of either the SegFormer, EfficientViT or U-Net (instead of increasing this volume by $+33\%$ as is common with RGB data). We claim that this informational bottleneck forces the model to get rid of most of the potentially valuable information right at the entry point to the model, which means that the later layers cannot effectively extract high level compressed information, as this extraction has to occur in the first convolutional layer.
@@ III-B Proposed architecture
Figure 5: Illustration of the Hyper SegFormer model for semantic segmentation of methane plumes in Hyperspectral data. We highlight in blue the three proposed modular adjustments: 1.) Spectral layers (denoted as “1x1 Conv”) in the Transformer blocks, 2.) Upscaling layer in the decoder network (denoted as “Upscale layer”) and finally 3.) adjusting the stride of the first Transformer block to 2 (denoted as “Stride”). In red we show an example of the progressive decimation of the resolution throughout the model (using all three adjustments).
@@ III-C Baselines
As a classical baseline approach for methane detection in hyperspectral data, we use the matched filter product (using the iterative variant mag1c [21]) with the threshold of 500 ppm×$\times$×m. Additionally, we apply an opening morphological filter to remove isolated noisy pixels in the product. As a classical machine learning baseline, we implement the HyperSTARCOP model based on the U-Net architecture proposed by [34] and train it on our new methane datasets. This model uses the pre-computed matched filter product alongside the RGB bands of the input data and effectively learns to clean the classical methane enhancement product. We note, that if the matched filter product misses a methane leak event, this model has no way of accessing the original information in the hyperspectral datacube to recover the detection.
@@ IV-A Hardware Restrictions
Therefore, with several tested model variants, we also conducted experiments measuring the runtime on compute constrained hardware, namely on the flatbed hardware provided by Unibap and D-Orbit [37]. These machines serve as a proxy to match the computational in-space environment of a D-Orbit’s ION-SCV004 satellite, which is equipped with: quad-core 1.2GHz, AMD GX-412HC SOC CPU processor, the Intel Movidius Myriad X VPU and 2GB of RAM. The same compute environment was used in [43] and similar environments leveraging Myriad chips were used on prior missions such as ESA’s PhiSat-1 [51] and PhiSat-2 [52].
We have also selected two other devices typically used for Internet-of-Things (IoT) applications, the CPU powered Raspberry Pi boards and the GPU equipped NVIDIA Jetson AGX Xavier board. As has been noted by NASA’s guidelines on using Raspberry Pi’s in space [53], they are not suggested for deployment in any mission critical tasks due to their lower robustness in extreme environments. They have been however used as a proxies for devices of similar compute power in [54]. We then use the more powerful NVIDIA Jetson AGX Xavier board equipped with an 8-core NVIDIA Carmel ARMv8.2 CPU at 2.265 GHz and a GV10B GPU based on the Volta architecture with 16GB memory (shared between CPU and GPU). We used both ONNX Runtime python libraries and the Tensor RT inference engine (using the trtexec command), as was also used in the work of [2]. While initial tests of the effects of radiation on NVIDIA’s Jetson devices was explored in [55] and proposed for space applications in [56], these works also state that a careful examination and future tests are still needed.
@@ V Results
Baseline, mag1c + morpho. N/A 37.65 31.69 46.37 23.19 89.0
@@ V-A Methane detection on EMIT
Table II shows the results of several tested model variants for both SegFormer and EfficientViT architectures compared with classical machine learning approach of [34] and baseline method using [21]. We note that all of the deep learning approaches significantly outperform the classical baseline. Furthermore, our proposed HyperSegFormer (ConvUpStride variant) gains over 43.85% in the IoU score (and 27.88% in F1 and 20.7% in AUPRC) over the HyperSTARCOP MF+RGB model. Our model also achieves over 155% in IoU score (and 96.9% in F1 score) over the classical matched filter baseline. These improvements hold over all measured metrics. We see increasing scores when implementing proposed adaptations in both the SegFormer (from 60.26 F1 score to 74.27) and EfficientViT (from 68.40 F1 score to 72.26) of our HyperspectralViT models.
@@
Baseline, mag1c + morpho. N/A 40.14 67.50 39.95 30.57 58.42 25.11 75.43
@@
Interestingly, when evaluating the models pre-trained on the dataset with synthetic events on a new dataset of real methane leak events collected by EMIT, we observe that the zero-shot generalisation performance is low (only two model variants reach the MF baseline). We however see that with fine-tuning for 10 epochs, we are able significantly outperform both the classical baselines and all model variants trained from scratch. Namely the HyperSegFormer (ConvUpStride variant) improves the performance of the MF baseline by 12.9% in F1 score (and 17.1% in IoU) and the same version of the model trained from scratch by over 6.9% in F1 score (and 8.8% in IoU). The HyperEfficientViT (ConvUp variant) outperforms the MF baselines by more than 23% in F1 (31.5% in IoU) and the same model variant trained from scratch by 19.2% in F1 (and 25.6% in IoU). Interestingly, the ConvUp variant slightly outperforms the ConvUpStride variants on this dataset with the EfficientViT models. Overall, we see, that on the task of detecting real methane leak events in EMIT data, we benefit from pre-training our models on a dataset of synthetically created events (especially in cases similar to ours, where the dataset of real events is relatively small).
@@ V-B Methane detection on AVIRIS-NG
When training our proposed models on the previously created benchmark dataset of real methane leak events in AVIRIS-NG data, we see some interesting mixed outcomes. While the performance of our models drops on F1 score of strong methane leak events (plume emissions larger than 1000 kg/h) in contrast to the HyperSTARCOP “MF+RGB” model, on the metrics that aggregate all event sizes together, we see improvements. Namely, the HypeSegFormer (ConvUpStride variant) outperforms the HyperSTARCOP “MF+RGB” baseline by 12.9% in F1 score and by 18% in IoU score. The threshold-less AUPRC metric is also improved by over 17%. The high performance of previous STARCOP models on strong methane leak events could be explained by the fact that the used matched filter method, mag1c, works better for large plumes (as it is iteratively increasing the contrast of positive pixels - this could however lead to lowered performance for weaker events).

### v3

@@ Abstract
On-board processing of hyperspectral data with machine learning models would enable an unprecedented amount of autonomy across a wide range of tasks allowing new capabilities such as early warning systems and automated scheduling across constellations of satellites. However, current classical methods suffer from high false positive rates and therefore prevent easy automation while previously published deep learning models exhibit prohibitive computational requirements. We propose fast and accurate machine learning architectures which support end-to-end processing of data with high spectral dimension without relying on hand-crafted products or spectral band compression techniques. We create three new large datasets of hyperspectral data containing all relevant spectral bands from the near global sensor EMIT. We evaluate our models on two tasks related to hyperspectral data processing - methane detection and mineral identification. Our models reach a new state-of-the-art performance on the task of methane detection, where we improve the F1 score of previous deep learning models by 27% on a newly created synthetic dataset and by 13% on the previously released large benchmark dataset. Our models generalise from synthetic datasets to data with real methane leak events and boost performance by 6.9% in F1 score in contrast with training models from scratch on the real data. Finally, with our newly proposed architectures, one capture from the EMIT sensor can be processed within 30 seconds on a realistic proxy of the ION-SCV 004 satellite and in less than 0.64 seconds on a GPU powered Jetson AGX Xavier board.
@@
publicationid: pubid: © 2025 The Authors. This work is licensed under a Creative Commons Attribution 4.0 License. For more information, see https://creativecommons.org/licenses/by/4.0/
@@ I Introduction
Mineral identification is a task related to scene decomposition into individual endmembers from a spectral library. On its own it could serve as a strong general task highly relevant for future Foundation models working with hyperspectral data. Hyperspectral Foundation models have been explored in the very recent works of [4, 5]. We note that one of the example downstream tasks of interest could also be methane detection, as it has been shown that having information about scene background is beneficial for separation of methane leak events from confounders [6].
@@
Figure 1: Illustration of the limitations of previous approaches. First, dependence on classically used products computed from hyperspectral data (such as matched filters in [17]) may lead to loss of accuracy due to imperfect capture of the events of interest. Second, off-the-shelf versions of machine learning models aren’t adapted to hyperspectral data and may lose valuable information in the early layers of the models. Finally, our adapted model leverages information from all relevant hyperspectral bands, which leads to improvements in both accuracy and inference speeds of our method.
@@ I-A Literature review
As has been identified by [32] and several overview papers [33, 34, 35], research in machine learning for processing hyperspectral data has been limited mainly due to the lack of large, publicly available datasets with high quality annotations. Many works still use very small datasets such as Indian Pines [36], University of Pavia, Houston or Salinas, all of which contain only a few data captures. As further noted by [33], this data is often divided into almost equal train and test subsets, which leads to model overfitting and results with very high accuracies almost regardless of the used model. In the domain of methane leak detection, only the recent works of [17, 6] release datasets created from a larger archive of full hyperspectral data captures provided by NASA. The work of [17] furthermore refines the available labels. The works of [37, 38] propose machine learning models for detection and quantification of methane leak events in hyperspectral data, however the used datasets and code/model weights are not published. In contrast, we publish all our used code and trained models and present the created datasets in a machine learning ready format.
@@ II-B Mineral aggregation
To create our new mineral dataset, we adapt the steps described in the Theoretical basis document [48] released alongside the “EMITL2BMIN.001” product [49] computed from the EMIT data. This data contains identifications of 381 individual components (separated into two groups). These can be aggreggated into larger mineral classes, for this we select the so-called “EMIT10” subset defined in [48]. The aggregation process is shown on Figure 4. We note that we use only the binary maps of three selected mineral classes out of the total 10. We chose the subset M = {Goethite, Hematite, Kaolinite}, as these classes were abundant in the locations sampled in our dataset. We use the released annotations which were computed using the Tetracorder method [27], and as such we can’t really use Tetracorder as a classical baseline (as it would unfairly obtain perfect scores). Therefore, we consider the created dataset as having pseudo-ground truth labels, and our results only as initial proof-of-principle. We however note, that given any manually validated dataset, our models could be easily retrained with the new data.
@@ II-C Final created datasets
Firstly, we created the “OxHyperSyntheticCH4” dataset of synthetically simulated methane leak events using the labels of the STARCOP dataset with newly downloaded near-global hyperspectral datacubes from the EMIT sensor (in L1B level of data processing). In total we keep 796 large tiles (of 512x512 px) in the training dataset, 198 in the validation dataset and 200 in the test dataset. Each subset is made from completely non-overlapping sources of data (clean EMIT datacubes and labels from the STARCOP dataset from AVIRIS-NG data). In each subset we uniformly sample the available plume sizes, so that we have similar distribution of event sizes. Half of these tiles contain simulated events while the other half is kept event free to provide negative samples. We note that these large tiles (of 512x512 px) are further tiled when training the models. All datasets created in this paper use tile size of 64 pixels with an overlap of 32 pixels. When evaluating the models (either on the Val or Test subsets), we keep the full resolution of 512x512 px. The synthetic Train dataset gets tiled into 174341 samples. The total size of this data is 228 GB.
Secondly, we create the “OxHyperRealCH4” dataset of real methane leak events. In total we have 279 large tiles in the training dataset, 91 in the validation dataset and 98 in the test dataset. When further tiled, the training dataset has 54805 samples. These have reliable ground truth labels, which we calculated from the original labels provided on the EMIT data portal and consequentially manually checked. We note that this dataset is smaller than the other ones and training a machine learning model from scratch could lead to overfitting and poor generalisation, however it can also be used for model fine-tuning and to test generalisation ability of models trained on other datasets. The total size of this data is 47 GB.
Thirdly, we create “OxHyperMinerals”, a completely new dataset for mineral identification, with labels of 3 selected mineral classes and source annotation for all 381 constituents (separated into two groups). In total we have 796 large tiles in the training dataset, 198 in the validation and finally 200 in the test dataset. When further tiled, the training dataset has 174341 samples. The large tiles exactly match the tiles used in the synthetic methane events dataset, however they do not contain the simulated events and contain more bands. The mineral dataset contains all 285 spectral bands in between spectral range of 381-2492nm. The total size of this data is 372 GB.
Finally, we also use a variant of the STARCOP dataset from [17] with all relevant spectral bands extracted from the source data. The STARCOP dataset uses tile size of 128 pixels with an overlap of 64 pixels for its Train subset. The Test subset is kept in its original resolution of 512x512 px. We train our models and compare them on this benchmark dataset with the previously reported results.
@@ III-A Limitations of existing approaches
We are interested in training these architectures in an end-to-end manner on the full ranges of selected bands of hyperspectral datacubes. In the methane detection task, this allows us to not depend on the classical matched filter product, which can be slow and which contains false detections of confounder materials. Models designed with typical computer vision data in mind (with one to three bands) however suffer from what we call informational bottleneck when they are used with data with large spectral dimensionality. As an example, if we select 86 bands of data which are relevant for our task of methane detection, the first layer in the vanilla versions of the used models reduces the data volume by $-95.34\%$ in the case of either the SegFormer, EfficientViT or U-Net (instead of increasing this volume by $+33\%$ as is common with RGB data). We claim that this informational bottleneck forces the model to get rid of most of the potentially valuable information right at the entry point to the model, which means that the later layers cannot effectively extract high level compressed information, as this extraction has to occur in the first convolutional layer.
@@ III-B Proposed architecture
Figure 5: Illustration of the Hyper SegFormer model for semantic segmentation of methane plumes in Hyperspectral data. We highlight in blue the three proposed modular adjustments: 1.) Spectral layers (denoted as “1x1 Conv”) in the Transformer blocks, 2.) Upscaling layer in the decoder network (denoted as “Upscale layer”) and finally 3.) adjusting the stride of the first Transformer block to 2 (denoted as “Stride”). In red we show an example of the progressive decimation of the resolution throughout the model (using all three adjustments) - we highlight that the typical off-the-shelf variants of the models reduce the output resolution by a factor of 4.
@@ III-C Baselines
As a classical baseline approach for methane detection in hyperspectral data, we use the matched filter product (using the iterative variant mag1c [24]) with the threshold of 500 ppm×$\times$×m. Additionally, we apply an opening morphological filter to remove isolated noisy pixels in the product. More specifically, we perform the erosion and dilation operations (using the implementations from the Kornia library). We explore two kernel shapes: “cross” kernel (3x3 pixels matrix with the middle row and column full of ones) and “ones” kernel (3x3 pixels matrix with all values set to ones). We observed small differences when trying these kernels, the “cross” kernel further corresponds to the method used in [17]. We observed that other kernel shapes, larger sizes, or different initial thresholds didn’t improve the results.
As a classical machine learning baseline, we implement the HyperSTARCOP model based on the U-Net architecture proposed by [17] and train it on our new methane datasets. This model uses the pre-computed matched filter product alongside the RGB bands of the input data and effectively learns to clean the classical methane enhancement product. We note, that if the matched filter product misses a methane leak event, this model has no way of accessing the original information in the hyperspectral datacube to recover the detection.
@@ IV-A Hardware Restrictions
Therefore, with several tested model variants, we also conducted experiments measuring the runtime on compute constrained hardware, namely on the flatbed hardware provided by Unibap and D-Orbit. These machines serve as a proxy to match the computational in-space environment of a D-Orbit’s ION-SCV004 satellite, which is equipped with: quad-core 1.2GHz, AMD GX-412HC SOC CPU processor, the Intel Movidius Myriad X VPU and 2GB of RAM. The power consumption of the used environment (SpaceCloud® iX5-100), is estimated to about 18W by [39]. The same compute environment was used in [45] and similar environments leveraging Myriad chips were used on prior missions such as ESA’s Phi-Sat-1 [53] and Phi-Sat-2 [54].
We have also selected two other devices typically used for Internet-of-Things (IoT) applications, the CPU powered Raspberry Pi boards and the GPU equipped NVIDIA Jetson AGX Xavier board. As has been noted by NASA’s guidelines on using Raspberry Pi’s in space [55], they are not suggested for deployment in any mission critical tasks due to their lower robustness in extreme environments. They have been however used as a proxies for devices of similar compute power in [56]. The estimated power consumption of Raspberry Pi devices is around 5W during a benchmark. We then use the more powerful NVIDIA Jetson AGX Xavier board equipped with an 8-core NVIDIA Carmel ARMv8.2 CPU at 2.265 GHz and a GV10B GPU based on the Volta architecture with 16GB memory (shared between CPU and GPU). The power consumption of this board can be adjusted, in our case we used the 30W mode. We used both ONNX Runtime python libraries and the Tensor RT inference engine (using the trtexec command), as was also used in the work of [2]. While initial tests of the effects of radiation on NVIDIA’s Jetson devices was explored in [57] and proposed for space applications in [58], these works also state that a careful examination and future tests are still needed.
@@ V Results
Baseline, mag1c, morpho. (cross 3x3) N/A 37.65 31.69 46.37 23.19 89.0
Baseline, mag1c, morpho. (ones 3x3) N/A 41.46 41.68 41.25 26.15 79.0
@@ V-A Methane detection on EMIT
Table II shows the results of several tested model variants for both SegFormer and EfficientViT architectures compared with classical machine learning approach of [17] and baseline method using [24] (regardless of the choice of kernel used in the baselines, although the “ones” baseline variant improves most of the scores over “cross” originally used in [17]). We note that all of the deep learning approaches significantly outperform the classical baseline. Furthermore, our proposed HyperSegFormer (ConvUpStride variant) gains over 43.85% in the IoU score (and 27.88% in F1 and 20.7% in AUPRC) over the HyperSTARCOP MF+RGB model. Our model also achieves over 155% in IoU score (and 96.9% in F1 score) over the classical matched filter baseline. These improvements hold over all measured metrics. We see increasing scores when implementing proposed adaptations in both the SegFormer (from 60.26 F1 score to 74.27) and EfficientViT (from 68.40 F1 score to 72.26) of our HyperspectralViT models.
@@
Baseline, mag1c, morpho. (cross 3x3) N/A 40.14 67.50 39.95 30.57 58.42 25.11 75.43
Baseline, mag1c, morpho. (ones 3x3) N/A 45.01 64.93 43.06 40.32 50.94 29.04 57.71
@@
Interestingly, when evaluating the models pre-trained on the dataset with synthetic events on a new dataset of real methane leak events collected by EMIT, we observe that the zero-shot generalisation performance is low (only two model variants reach the MF baseline - the “cross” kernel variant). We however see that with fine-tuning for 10 epochs, we are able significantly outperform both the classical baselines and all model variants trained from scratch. Namely the HyperSegFormer (ConvUpStride variant) improves the performance of the MF baseline by 12.9% in F1 score (and 17.1% in IoU) and the same version of the model trained from scratch by over 6.9% in F1 score (and 8.8% in IoU). The HyperEfficientViT (ConvUp variant) outperforms the MF baselines by more than 23% in F1 (31.5% in IoU) and the same model variant trained from scratch by 19.2% in F1 (and 25.6% in IoU). Interestingly, the ConvUp variant slightly outperforms the ConvUpStride variants on this dataset with the EfficientViT models. Overall, we see, that on the task of detecting real methane leak events in EMIT data, we benefit from pre-training our models on a dataset of synthetically created events (especially in cases similar to ours, where the dataset of real events is relatively small).
@@ V-B Methane detection on AVIRIS-NG
When training our proposed models on the previously created benchmark dataset of real methane leak events in AVIRIS-NG data, we see some interesting mixed outcomes. While the performance of our models drops on F1 score of strong methane leak events (plume events larger than 1000 px) in contrast to the HyperSTARCOP “MF+RGB” model, on the metrics that aggregate all event sizes together, we see improvements. Namely, the HypeSegFormer (ConvUpStride variant) outperforms the HyperSTARCOP “MF+RGB” baseline by 12.9% in F1 score and by 18% in IoU score. The threshold-less AUPRC metric is also improved by over 17%. The high performance of previous STARCOP models on strong methane leak events could be explained by the fact that the used matched filter method, mag1c, works better for large plumes (as it is iteratively increasing the contrast of positive pixels - this could however lead to lowered performance for weaker events).